OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the third measure is the hypotenuse. 6, 12, 18 Answer A. yes B. no

OpenStudy (amistre64):

whats the pythag thrm state?

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (binary3i):

no

OpenStudy (amistre64):

good, then all we have to do is use these numbers

OpenStudy (amistre64):

6^2 + 12^2 =? 18^2

OpenStudy (anonymous):

ooo (: i just needed a start (: thankss

OpenStudy (anonymous):

look the pythagora's theorom is a^2+b^2 = c^2 Base^2+Perpendicular^2=Hypotenuse^2 18^2= 12^2 + 6^2 check if this is right!